OpenStudy (anonymous):

Need to find f'(0) of this function

OpenStudy (anonymous):

\[\frac{ e^x g(x) }{ h(x)+2g(x) }\]

OpenStudy (anonymous):

g(0) is 2, g'(0) is 5, h(0) is 1, h'(0) is -1

OpenStudy (anonymous):

do u know the quotient rule?

OpenStudy (anonymous):

oh im dumb plug it in of g(x) with x as 0, and then derive the resulting thing, (its very easy)

OpenStudy (amistre64):

\[\frac{ e^x g(x) }{ h(x)+2g(x) }\] \[ e^x~ g~ ({ h+2g })^{-1}\] \[ e^x~ g~ ({ h+2g })^{-1}\\ +e^x~ g'~ ({ h+2g })^{-1}\\ +e^x~ g~ ({ h+2g })'^{-1}\] \[({ h+2g })'^{-1}=-({ h+2g })^{-2}~( h'+2g')\]

OpenStudy (amistre64):

quotient rule is overrated, product rule is much simpler to keep track of

OpenStudy (amistre64):

x = 0 g=2, g'=5 h=1, h'=-1 plug and play

OpenStudy (anonymous):

cant u differentiate e^x * 2 / (1+2*2) ?? @amistre64

OpenStudy (anonymous):

and then u get 2/5 e^x as ur derivative

OpenStudy (amistre64):

well, the functions are not constant in their definitions, but lets see what out final result is

OpenStudy (anonymous):

yah but its only the derivative at 0 so?

OpenStudy (amistre64):

\[ 2/~ ({ 1+2*2 })\\ + 5/~ ({ 1+2*2 })\\ - 2~ (-1+2*5)/({ 1+2*2 })'^{2}\] \[ 2/5+ 5/5 - 18/25\]

OpenStudy (amistre64):

so, 7/5 - 18/25 (35- 18)/25 17/25 not= 2/5

OpenStudy (amistre64):

so no, this is not the same as the derivative of 2/5 e^x

OpenStudy (amistre64):

in otherwords, the function at x=0 is 2/5, but the slope of the line at x=0 is 17/25 assuming of course i didnt make any errors